# On the Convergence Rate of Linear $\text{\sf Datalog}^{\circ}$ over Stable Semirings

### v1

@@ 2 Preliminaries
[…]
@@
Proof.
This follows directly from the $p$-stability of 1, and the fact that $11^{2}=$ 1. ∎
Let us now explain the general procedure for creating a matrix $A$ and a vector $b$ from a linear $\text{\sf Datalog}^{\circ}$ program $Q$. Each ground tuple for each IDB predicate $R$ can be viewed as a variable, and each ground tuples of EDB predicates can be viewed as constants. Then a $\text{\sf Datalog}^{\circ}$ rule, where the head is IDB predicate $R$, can be converted into a collection of linear equations, one for each ground tuble of $R$. Since all rules that share IDB $R$ as a the head can be combined via $\oplus$, we can compactly rewrite the entire set of $\text{\sf Datalog}^{\circ}$ rules as a collection of linear equations of the following form:
@@ 3 Upper bounding the Convergence as a Function of the Matrix Dimension and the Semiring Stability
In Lemma 3.3 we show that a particular statement about $W$ is invariant through the recursive construction. A proof Lemma 3.3 requires the following lemma. The proof of the following lemma (or at least, the techniques needed for a proof) can be found in most introductory graph theory texts, e.g. [27].
@@
Proof.
The proof is by induction on the number of steps of the recursive construction. The statement is obviously true for the initial walk $W$, which is the base case. Now consider one step of the recursive construction. Removing copies of a cycle from $W$ does not change the difference between the in-degree and out-degree of any vertex. Thus by Lemma 3.2 the only issue we need to consider is vertex $i$ and vertex $j$ possibly ending up in different connected components of $W^{\prime}$. Let $D_{f}$ be the connected component of $W$ that contains $i$ (which also contains $j$ by the induction hypothesis), let $D^{\prime}_{a}$ be the connected component of $W^{\prime}$ that contains $i$, and let $D^{\prime}_{b}$ be the connected component of $W^{\prime}$ that contains $j$, where $a\neq b$. Then the walk in $D_{f}$ from $i$ to $j$ must cross the cut (in either direction) formed by the vertices in $D^{\prime}_{a}$ an odd number of times. But the edges in $C_{h}$ must cross the cut (in either direction) formed by the vertices in $D^{\prime}_{a}$ an even number of times. Thus there must be an edge in $D_{f}$ minus $z_{h}$ copies of $C_{h}$ that must cross the cut (in either direction) formed by the vertices in $D^{\prime}_{a}$. However, then this is a contradiction to $D^{\prime}_{a}$ be a connected component in $W^{\prime}$. ∎
@@
Proof.
As $W$ contains $k+1=n(n^{2}-n)(p+2)+n$ edges, then it must contain a simple cycle, which implies $\ell\geq 1$. Consider one iteration of our recursive construction. There must be a directed edge $e\in C_{h}$ that appears exactly $z_{h}$ times in $W$. Thus there are no occurrences of $e$ in $W^{\prime}$. Thus $e$ can not appear in any future cycles, that is $e\notin C_{g}$ for any $g>h$. The first observation then follows because there are at most $n^{2}-n$ different edges in $G$. The second observation is then an immediate consequence of the first observation, and the invariant established in Lemma 3.3 . The third observation follows because no edges are ever lost or created in the recursive construction. ∎
Observation 2.
@@
Proof.
Since $P$ is a simple path it contains at most $n-1$ edges. Thus $W-P$ contains at least $n(n^{2}-n)(p+2)$ edges. As any simple cycle contains at most $n$ edges, and as there are at most $\ell\leq n^{2}-n$ cycles, then by applying the pigeon hole principle to the cycle decomposition of $W$ we can conclude that there must be a cycle $C_{s}$ that has multiplicity at least $p+2$, that is $z_{s}\geq p+2$. ∎
For convenience, consider renumbering the cycles so that $s=1$ where $z_{s}\geq p+1$, which exists by Observation 2. For each $h$ such that $11\leq h\leq p+$ 1, let define $W_{h}$ be the collection of edges in $W$ minus $h$ copies of every edge in $C_{1}$.
Observation 3.
@@
Proof.
As $z_{1}\geq p+2$ and $h\leq p+1$, every edge that appears in $W$ also appears in $W_{h}$. So $W_{h}$ has the same connectivity properties as $W$, and $W_{h}$ has the same vertices with positive in-degree as does $W$. Also as $C_{1}$ is a simple cycle, the difference between in-degree and out-degree for each vertex is the same in $W_{h}$ as in $W$. Thus the result follows by appealing to Lemma 3.2. ∎
Observation 4.
@@
Proof.
This follows directly from the defintion of $W_{h}$. ∎
Observation 5.
@@
Proof.
This follows because $W$ has $k+1$ edges and $C_{1}$ is non-empty, so removing edges in $C_{i}$ strictly decreases the number of edges. ∎
We are now ready to prove Lemma 3.1. By Observation 5 we know that each $\Phi(W_{h})$ is included in $A^{(k)}$, and thus there exists an element $r$ in the semiring $S$ such that
@@
The equality in line (7) follows from Observation 4. The equality in line (8) follows from the defintion of $\Phi$. The key step in this line of equations is the equality in line (9), which follows from the stability of $\Phi(C_{1})$. The rest of the equalities follow from basic algebraic properties of semirings, or by definition of the relavent term.
@@
4 A Lower Bound for Convergence Rate of Linear $\text{\sf Datalog}^{\circ}$ Programs
This section lower bounds the convergence rate of linear $\text{\sf Datalog}^{\circ}$ programs using naive evaluation and establishes Theorem  1.7. In particular, this section will construct a semiring and a datalog program that requires $\Omega(pn^{3})$ iterations to converge. The section first constructs a semiring then defines a matrix $A$ and finally the converge rate is bounded.
4.1 Constructing the Semiring
The ground set $S$ consists of multi-sets of $\{1,\ldots m\}$; later we will set $m=\Theta(n^{2})$. To avoid confusion, we will refer to $1,2,\ldots m$ as items, which will be distinguished from the elements of $S$. An element in $S$ can have up to $p\geq 1$ copies of each item $i\in[m]$. Additionally, there is a special element $\mathcal{O}$ in the semiring. Thus, the ground set $S$ has $(p+1)^{m}+1$ elements.
For a multiset $A$ we denote the number of copies of item $i$ in $A$ as $A_{i}$. We now define the semiring operations. Consider two distinct elements (multisets) $A$ and $B$ in $S\setminus\{\mathcal{O}\}$. Define $C=A+B$, where $C_{i}=\max\{A_{i},B_{i}\}$ for all $i\in[m]$. Further, $\mathcal{O}+A=A+\mathcal{O}=A$ for all $A\in S$. Multiplication $C=A\cdot B$ is defined as follows: $C_{i}=\min\{A_{i}+B_{i},p\}$ for all $i\in[m]$. Further $\mathcal{O}\cdot A=A\cdot\mathcal{O}=\mathcal{O}$ for all $A\in S$.
We establish that this is a semiring where $\mathcal{O}$ is the additive identity and the empty multi-set $\emptyset$ is the multiplicative identity (i.e. $1$). By definition, $\mathcal{O}$ is the annihilator element. We show in Appendix B that this is a commutative semiring. The proof easily follows from the definitoin of the semiring.
4.2 Defining the Matrix $A$
@@
This section defines the matrix $A$. To do so, we first define a graph $G$. Let $G$ be a directed graph where the vertex set is the integers from 1 to $n$ inclusive. For simplicity assume $n$ is divisible by 3. The vertices are partitioned into 3 parts, $B=\{1,\ldots,n/3\}$, $C=\{n/3+1,\ldots,2n/3\}$ and $D=\{2n/3+1,\ldots n\}$. There is a directed edge from each vertex in $B$ to vertex $n/3+1$, there is a directed edge from vertex $32n/$ 3 to each vertex in $D$, and there is a directed edge from each vertex in $D$ to each vertex in $B$. Finally, all vertices in $C$ are sequentially connected from $n/3+1$ to $32n/$ 3, i.e., there is a directed edge from $\tau$ to $\tau+1$ for all $\tau\in[n/3+1,2n/3-1]$.
Index the edges by $1$ through $m$ and assign distinct labels (items) to them. So, $m$ is exactly equal to the number of distinct items. Notice that $m$ is $\Theta(n^{2})$. If there is a directed edge $(i,j)$ with label $k$ in $G$ then $A_{i,j}=\{k\}$. This is the element corresponding to the multiset with one copy of $k$. If there is no directed edge $(i,j)$ in $G$ then $A_{i,j}=\mathcal{O}$.
Lemma 4.1.
The number of steps until convergence is $\Omega(pn^{3})$.
Proof.
@@
Note that there are $|B|\cdot|D|=n^{2}/9$ edges from $D$ to $B$. Consider a long walk, say from $i:=n/3+1$ to $j:=n/3+2$. Observe that the walk must visit all edges within $C$ (and the edge from $32n/$ 3 to $12n/3+$ 1) before visiting exactly one edge from $D$ to $B$. Thus, to visit each edge from $D$ to $B$ at least $p$ times, the walk must have length at least $p|B|\cdot|D|\cdot|C|=pn^{3}/27$. Similarly, there is such a walk of length at most $p|B|\cdot|D|(|C|+2)+1\leq 4pn^{3}$. Thus, we have shown that $A^{(k)}$ includes the multiset $Q$ that has $p$ copies of each item when $k\geq 4pn^{3}$. Further, we have shown that $A^{(k)}$ doesn’t include $Q$ when $k<p(n/3)^{3}$. This proves the lower bound on the convergence rate. ∎
@@ 5 Bounding Convergence in Terms of the Semiring Ground Set Size
where $q\geq p+1$. Here $\mathcal{C}^{\prime}\setminus C^{q}$ implies the resulting collection of cycles we obtain after removing $q$ copies of $C$ from $\mathcal{C}^{\prime}$. Now consider walk $W_{q^{\prime}}$ that concatenates $T$, $q^{\prime}$ copies of $C$, $\mathcal{C}^{\prime}\setminus C^{q^{\prime}}$, and $T^{\prime}$. Here, the walk starts with $T$ and ends with $T^{\prime}$, and the cycles can be placed in an arbitrary order. This is because $T$ is a walk from $i$ to $v$, and all the cycles in $\mathcal{C}^{\prime}$ start from $v$ and end at $v$—due to the fifth property—and $T^{\prime}$ is a walk from $v$ to $j$ Further, they are all shorter than $W$ because $W_{q}$ is no longer than $W$ due to the second property. Therefore, $W_{0},W_{1},\ldots W_{q-1}\in\mathcal{W}_{i,j}^{(k)}$. Thus, thanks to $p$-stability, we have $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$ as desired. ∎
Although we gave the full proof of Theorem 1.8, to convey intuition better, we also give some warm-up analyses in Appendix C by giving a looser bound for the general case and subsequently by considering a special case of $p=1$.
6 Lower Bounds on Convergence in Terms of the Semiring Ground Set Size
This section constructs a lower bound of the convergence rate in terms of the size of the ground set. The goal is to show Theorem 1.9, which is implied by the following lemma.
Lemma 6.1.
There exists an idempotent semiring on $L$ elements and a matrix $A$ of size $n$ by $n$ that requires $\Omega(nL)$ steps to converge to a fixed point.
Proof.
Consider a semiring that whose all powers of 2 from $1$ to $2^{L}$ and the value 00. The value of $2^{L}$ is the largest value. Summation $A$ and $B$ in the semiring is $\min\{A\cdot B,2^{L}\}$, where $\cdot$ is standard multiplication. Addition is standard maximum. By definition, addition is idempotent ensuring the semiring is idempotent.
The matrix $A$ corresponds to a computation graph with a cycle of length $n$. All edges that are not in cycles are labeled 00. All edges of the cycle are labeled $1$, the identity in standard multiplication, except for one edge, which is $2$. Notice that multiplying all edges of the cycle $i$ times results in the symbol $2^{i}$.
The cycle needs to be traversed $L$ times to reach $2^{L}$. The walk is of length $\Theta(nL)$.∎
Lemma 6.2.
There exists a matrix $A$ of size $n$ by $n$ which requires $\Omega(np\frac{\log L}{\log p})$ steps to find a fixed point over a semiring of $L$ elements that is $p$-stable.
Proof.
Consider the following semiring. The $L$ elements are over vectors of $\ell$ dimensions. Each position can be $0,1,2,\ldots,p$. Consider any two elements $x$ and $y$. Let $x_{i}$ and $y_{i}$ be the $i$th dimension of $x$ and $y$, respectfully. The addition operation on $x$ and $y$ returns whichever among $x$ and $y$ are lexicographically larger. Multiplication of $x$ and $y$ produces the vector $z$ where $z_{i}=\min\{x_{i}+y_{i},p\}$.
We first claim that this is a semiring. Notice that the all 00 vector is a monoid for multiplication. The vector of all $p$’s is the multiplicative identity. The only case that is non-obvious is that multiplication distributes over addition. Consider three vectors $a$, $b$ and $c$. Consider the expression $c\cdot(a+b)$. We aim to show that this is equal to $c\cdot a+c\cdot b$. Without loss of generality say $a$ is lexicographically bigger than $b$. Then we have that $c\cdot(a+b)=c\cdot a$ because $a+b=a$ using that $a$ is lexiographically bigger than $c$. Similarly, $c\cdot a+c\cdot b=c\cdot a$ because multiplication is standard addition and $a$ is lexicographically bigger than $b$.
The matrix $A$ corresponds to the following graph. There are two special nodes $a$ and $b$. They are connected by $\ell$ one-hop path using two edges from $a$ to $b$. The $k$th path goes via a node $c_{k}$. The edge from $a$ to $c_{k}$ is labeled with the 00 vector. The edge from $c_{k}$ to $b$ is labeled with a vector of all 00s except a 1 in dimension $k$. Additionally, $b$ is connected to $a$ via a directed path of length $n-\ell-2$. The edges of this path are all labeled with the all 00 vector.
To collect the vector consisting of $p$ in each dimension, one needs to walk a cycle starting at $a$ at least $p\ell$ times. To see why, notice multiplication of the edges corresponding to a single time-around cycle increases a single dimension by at most one. Moreover, addition’s definition ensures the final output is the vector that is lexicographically the biggest among each walk. Each cycle is of length $\Omega(n)$. The length of the walk required is $\Omega(p\ell n)$. Setting $\ell=\frac{\log L}{\log p}$ gives the lemma by noting that $L=p^{\ell}$ is the number of elements.∎

### v2

@@ 2 Preliminaries
$\bm{x}\otimes_{p}\bm{y}\stackrel{{\scriptstyle\text{def}}}{{=}}$ ${\min}_{p}(\bm{x}+\bm{y})$ $\bm{1}_{p}\stackrel{{\scriptstyle\text{def}}}{{=}}$$\{0,\infty,\ldots,\infty\}$
@@
Proof 2.3.
This follows directly from the $p$-stability of 1, and the fact that $11^{2}=$ 1.
Let us now explain the general procedure for creating a matrix $A$ and a vector $b$ from a linear $\text{\sf Datalog}^{\circ}$ program $Q$. Each ground tuple for each IDB predicate $R$ can be viewed as a variable, and ground tuples of EDB predicates can be viewed as constants. Then a $\text{\sf Datalog}^{\circ}$ rule, where the head is IDB predicate $R$, can be converted into a collection of linear equations, one for each ground tuble of $R$. Since all rules that share IDB $R$ as a the head can be combined via $\oplus$, we can compactly rewrite the entire set of $\text{\sf Datalog}^{\circ}$ rules as a collection of linear equations of the following form:
@@ 3 Upper bounding the Convergence as a Function of the Matrix Dimension and the Semiring Stability
In Lemma 3.3 we show that a particular statement about $W$ is invariant through the recursive construction. A proof Lemma 3.3 requires the following lemma. The proof of the following lemma (or at least, the techniques needed for a proof) can be found in most introductory graph theory texts, e.g. [22] Theorem 23.1.
@@
Proof 3.4.
The proof is by induction on the number of steps of the recursive construction. The statement is obviously true for the initial walk $W$, which is the base case. Now consider one step of the recursive construction. Removing copies of a cycle from $W$ does not change the difference between the in-degree and out-degree of any vertex. Thus by Lemma 3.2 the only issue we need to consider is vertex $i$ and vertex $j$ possibly ending up in different connected components of $W^{\prime}$. Let $D_{f}$ be the connected component of $W$ that contains $i$ (which also contains $j$ by the induction hypothesis), let $D^{\prime}_{a}$ be the connected component of $W^{\prime}$ that contains $i$, and let $D^{\prime}_{b}$ be the connected component of $W^{\prime}$ that contains $j$, where $a\neq b$. Then the walk in $D_{f}$ from $i$ to $j$ must cross the cut (in either direction) formed by the vertices in $D^{\prime}_{a}$ an odd number of times. But the edges in $C_{h}$ must cross the cut (in either direction) formed by the vertices in $D^{\prime}_{a}$ an even number of times. Thus there must be an edge in $D_{f}$ minus $z_{h}$ copies of $C_{h}$ that must cross the cut (in either direction) formed by the vertices in $D^{\prime}_{a}$. However, then this is a contradiction to $D^{\prime}_{a}$ be a connected component in $W^{\prime}$.
@@
Proof 3.5.
As $W$ contains $k+1=n(n^{2}-n)(p+2)+n$ edges, then it must contain a simple cycle, which implies $\ell\geq 1$. Consider one iteration of our recursive construction. There must be a directed edge $e\in C_{h}$ that appears exactly $z_{h}$ times in $W$. Thus there are no occurrences of $e$ in $W^{\prime}$. Thus $e$ can not appear in any future cycles, that is $e\notin C_{g}$ for any $g>h$. The first observation then follows because there are at most $n^{2}-n$ different edges in $G$. The second observation is then an immediate consequence of the first observation, and the invariant established in Lemma 3.3 . The third observation follows because no edges are ever lost or created in the recursive construction.
{observation}
@@
Proof 3.6.
Since $P$ is a simple path it contains at most $n-1$ edges. Thus $W-P$ contains at least $n(n^{2}-n)(p+2)$ edges. As any simple cycle contains at most $n$ edges, and as there are at most $\ell\leq n^{2}-n$ cycles, then by applying the pigeon hole principle to the cycle decomposition of $W$ we can conclude that there must be a cycle $C_{s}$ that has multiplicity at least $p+2$, that is $z_{s}\geq p+2$.
For convenience, consider renumbering the cycles so that $s=1$ where $z_{s}\geq p+1$, which exists by Observation 3. For each $h$ such that $11\leq h\leq p+$ 1, let define $W_{h}$ be the collection of edges in $W$ minus $h$ copies of every edge in $C_{1}$.
{observation}
@@
Proof 3.7.
As $z_{1}\geq p+2$ and $h\leq p+1$, every edge that appears in $W$ also appears in $W_{h}$. So $W_{h}$ has the same connectivity properties as $W$, and $W_{h}$ has the same vertices with positive in-degree as does $W$. Also as $C_{1}$ is a simple cycle, the difference between in-degree and out-degree for each vertex is the same in $W_{h}$ as in $W$. Thus the result follows by appealing to Lemma 3.2.
{observation}
@@
Proof 3.8.
This follows directly from the defintion of $W_{h}$.
{observation}
@@
Proof 3.9.
This follows because $W$ has $k+1$ edges and $C_{1}$ is non-empty, so removing edges in $C_{i}$ strictly decreases the number of edges.
We are now ready to prove Lemma 3.1. By Observation 3 we know that each $\Phi(W_{h})$ is included in $A^{(k)}$, and thus there exists an element $r$ in the semiring $S$ such that
@@
The equality in line (10) follows from Observation 3. The equality in line (11) follows from the defintion of $\Phi$. The key step in this line of equations is the equality in line (12), which follows from the stability of $\Phi(C_{1})$. The rest of the equalities follow from basic algebraic properties of semirings, or by definition of the relavent term.
@@
4 A Failed Lower Bound Attempt for Convergence Rate of Linear $\text{\sf Datalog}^{\circ}$ Programs
Recall that the definition of $p$-stability is that
$1\oplus u\oplus u^{2}\oplus\cdots\oplus u^{p}=1\oplus u\oplus u^{2}\oplus\cdots\oplus u^{p}\oplus u^{p+1}$
Thus it is natural (but wrong) to conjecture that in order for a walk $W$ to be absorbed by shorter walks that it must be the case that one can express $\Phi(W)$ as $XY^{p+1}$ where $X$ and $Y^{p+1}$ are products of two parts of some partition of $W$. In this case then there is at least the possibility that $W$ is absorbed by
$X\oplus X\otimes Y\oplus X\otimes Y^{2}\ldots\oplus X\otimes Y^{p}$
if it happened that there were shorter walks with products equal to these summands. We now demonstrate that in order to guarantee that a walk $W$ has the property that it can be decomposed so that $\Phi(W)=X\otimes Y^{p+1}$ then $W$ has to be of length $\Omega(pn^{3})$.
@@ 4 A Failed Lower Bound Attempt for Convergence Rate of Linear $\text{\sf Datalog}^{\circ}$ Programs
We define a graph $G$. Let $G$ be a directed graph where the vertex set is the integers from 1 to $n$ inclusive. For simplicity assume $n$ is divisible by 3. The vertices are partitioned into 3 parts, $B=\{1,\ldots,n/3\}$, $C=\{n/3+1,\ldots,2n/3\}$ and $D=\{2n/3+1,\ldots n\}$. There is a directed edge from each vertex in $B$ to vertex $n/3+1$, there is a directed edge from vertex $32n/$ 3 to each vertex in $D$, and there is a directed edge from each vertex in $D$ to each vertex in $B$. Finally, all vertices in $C$ are sequentially connected from $n/3+1$ to $32n/$ 3, i.e., there is a directed edge from $\tau$ to $\tau+1$ for all $\tau\in[n/3+1,2n/3-1]$. See Figure 1.
Figure 1: The graph $G$.
@@
Note that there are $|B|\cdot|D|=n^{2}/9$ edges from $D$ to $B$. Consider a long walk, say from $i:=n/3+1$ to $j:=n/3+2$, that visits each edge from $D$ to $B$ exactly $p$ times. Clearly, for such a walk $W$, we cannot express $W$ as $XY^{p+1}$ for some $X$ and $Y$. Observe that the walk must visit all edges within $C$ (and the edge from $32n/$ 3 to $12n/3+$ 1) before visiting exactly one edge from $D$ to $B$. Thus, to visit each edge from $D$ to $B$ at least $p$ times, the walk must have length at least $p|B|\cdot|D|\cdot|C|=pn^{3}/27$.
@@ 5 Bounding Convergence in Terms of the Semiring Ground Set Size
where $q\geq p+1$. Here $\mathcal{C}^{\prime}\setminus C^{q}$ implies the resulting collection of cycles we obtain after removing $q$ copies of $C$ from $\mathcal{C}^{\prime}$. Now consider walk $W_{q^{\prime}}$ that concatenates $T$, $q^{\prime}$ copies of $C$, $\mathcal{C}^{\prime}\setminus C^{q^{\prime}}$, and $T^{\prime}$. Here, the walk starts with $T$ and ends with $T^{\prime}$, and the cycles can be placed in an arbitrary order. This is because $T$ is a walk from $i$ to $v$, and all the cycles in $\mathcal{C}^{\prime}$ start from $v$ and end at $v$—due to the fifth property—and $T^{\prime}$ is a walk from $v$ to $j$ Further, they are all shorter than $W$ because $W_{q}$ is no longer than $W$ due to the second property. Therefore, $W_{0},W_{1},\ldots W_{q-1}\in\mathcal{W}_{i,j}^{(k)}$. Thus, thanks to $p$-stability, we have $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$ as desired.
Although we gave the full proof of Theorem 1.5, to convey intuition better, we also give some warm-up analyses in Appendix B by giving a looser bound for the general case and subsequently by considering a special case of $p=1$.